The charge density of a uniformly charged disk 0.410 m in diameter is 2.33 10-2 C/m2. What is the magnitude of the electric field along the disks axis at the following distances from its center? (a) 2.05 cm 121007x Your response is within 10% of the correct value. This may be due to roundoff error, or you could have a mistake in your calculation. Carry out all intermediate results to at least four-digit accuracy to minimize roundoff error. N/C (b) 20.5 cm (c) 41.0 cm N/C (d) 4.10 m N/C Need Help?Read H
Answer #1

For a disk of radius R and surface charge density sigma, the Electric field at a distance x on the axis of the disk is given by

EF

Therefore, Here R=D/2 = 0.410/2 =0.205 m, and charge density= 2.33 x 10^-2

Therefore for different valyes of x for different parts

2- -2 2 2 0.41 0.20SEF.AT x1 X10, 1 X 2.33X10 It- Ч.10